# Revisiting the predictability of dynamical systems: a new local data-driven approach

### v1

@@ Abstract
Nonlinear dynamical systems are ubiquitous in nature and they are hard to forecast. Not only they may be sensitive to small perturbations in their initial conditions, but they are often composed of processes acting at multiple scales. Classical approaches based on the Lyapunov spectrum rely on the knowledge of the dynamic forward operator, or of a data-derived approximation of it. This operator is typically unknown, or the data are too noisy to derive a faithful representation. Here, we propose a new data-driven approach to analyze the local predictability of dynamical systems. This method, based on the concept of recurrence, is closely linked to the well-established framework of local dynamical indices. Applied to both idealized systems and real-world datasets, this new index shows results consistent with existing knowledge, proving its effectiveness in estimating local predictability. Additionally, we discuss its relationship with local dynamical indices, illustrating how it complements the previous framework as a more direct measure of predictability. Furthermore, we explore its reflection of the scale-dependent nature of predictability, its extension that includes a weighting strategy, and its real-time application. We believe these aspects collectively demonstrate its potential as a powerful diagnostic tool for complex systems.
@@ 1 Introduction
The origin of this practical unpredictability may stem either from an intrinsic stochastic nature of the system or from the ignorance of the current state of the system and/or of the dynamic rule controlling its evolution [23] or from a combination of both. Furthermore, it is well-known that even low-dimensional nonlinear systems may be difficult to forecast if they exhibit an exponential divergence of close-by states [19]. In such cases the unpredictability comes from our lack of precise knowledge of the current state of the system: if we knew exactly the initial conditions we would always get the same future evolution. A common quantity used to characterize the predictability of a system is the maximum Lyapunov exponent $\lambda_{\rm{max}}$. If we look at the evolution of two trajectories starting from two points close to each other, namely ${\bf{x}}(t_{0})$ and ${\bf{x}}(t_{0})+{\delta}{\bf{x}}$, these will diverge exponentially fast as time progresses, until their distance will plateau around a value bounded by the diameter of the attractor [24]. Even more instructive is to look at the evolution of a ball around the state ${\bf{x}}(t)$. From this, one can define a full Lyapunov spectrum, made of Lyapunov exponents equal in number to the number of variables used to describe the phase space. The Lyapunov exponents are mean logarithmic growth rates of the lengths of the principal axes of this ball evolving under the action of the dynamic rule [25, 4, 24, 5]. The existence of at least one positive Lyapunov exponent is the hallmark of chaotic dynamical systems. Therefore, measuring predictability in dynamical systems can be re-framed as computing the divergence of a system’s trajectories over time.
@@
Besides Lyapunov exponents and their variants, information theory also offers methods to quantify predictability [39, 23]. Within this framework, various entropy-based indicators are usually employed as predictability measures for model-based prediction, with varying degree of success [40, 41, 42, 43]. However, because these methods are model-based, they are unsuitable for observational datasets. Other indices from nonlinear time series analysis also use predictability measures such as sample entropy, permutation entropy, and others [44, 45, 46]; nevertheless, they are only effective with time series or systems with relatively low dimensionality. To the best of the authors’ knowledge, no information theory method can infer local predictability information from high-dimensional observational data.
From the brief review of relevant literature just outlined, the above mentioned indices have several drawbacks. More specifically, Lyapunov exponents and their variants may not work well for nonlinear regimes, while information-theory-based methods may not be applicable to high-dimensional dynamical systems. In the next section, we leverage a recent framework proposed by Lucarini et al., [47], that blends extreme value theory [47] and the Poincaré recurrence theorem [2] to build a new local predictability approach.
@@
Lucarini et al., [47] introduced a novel framework to measure the local properties of dynamical systems. In particular, they combine extreme value theory [48] and the Poincaré recurrence theorem [2] to construct purely data-driven local dynamical indices, that can measure the local dimension of dynamical systems and their persistence at any point in time. Under the ergodic assumption, the Poincaré recurrence theorem guarantees that the system will visit an arbitrarily small neighborhood of any state again and again as time passes. In this methodology, for one given state of interest ${\boldsymbol{\zeta}}={\bf{x}}(t={t}_{\boldsymbol{\zeta}})$ in the phase space, its neighboring states, also known as recurrences, are used to characterize its local dynamical properties. To achieve this, we first define negative logarithmic returns as:
@@
where the dist function can be any distance metric. We note that requiring a trajectory falls within a neighborhood of ${\boldsymbol{\zeta}}$ (intended as a ball) is a synonym of having the time series $g[{\boldsymbol{\zeta}},{\bf{x}}(t)]$ above one threshold $s(q,{\boldsymbol{\zeta}})$, where $s(q,{\boldsymbol{\zeta}})$ is a high threshold associated to a quantile $q$ of the series $X\equiv g(x(t))$.
Then, we define a quantity called exceedance for the neighboring states (i.e., recurrences) of $\zeta$ as follows:
${\bf{u}}({\boldsymbol{\zeta}})=g({\boldsymbol{\zeta}},{\bf{x}}(t))-s(q,{\boldsymbol{\zeta}}),\;\;\;\;\forall~{}g({\boldsymbol{\zeta}},{\bf{x}}(t))>s(q,{\boldsymbol{\zeta}})$ (2)
According to extreme value theory [49, 50], assuming the independence of the exceedances, the cumulative probability distribution $F({\bf{u}},{\boldsymbol{\zeta}})$ follow the exponential member of the generalized Pareto distribution (GPD), that is:
The parameter $\sigma({\boldsymbol{\zeta}})$, which is the scale parameter of the distribution, depends on the state $\zeta$ in phase space and can be used to derive the local dimension via $d({\boldsymbol{\zeta}})=1/\sigma({\boldsymbol{\zeta}})$ [47]. In practice, the local dimension is computed by explicitly fitting an exponential distribution to the exceedances $\bf{u}({\boldsymbol{\zeta}})$, see discussion in [51] and [52]. The estimate of the attractor’s dimension derived from $d({\boldsymbol{\zeta}})$ seems to be more efficient than previous methods based on evaluating correlation integrals [53, 54]. The other local index, persistence ($\theta^{-1}$), is defined as the inverse of the extremal index $0\leq\theta\leq 1$ [55, 56], a dimensionless parameter that measures the inverse of the length of clustering of extremes. In this context, extremes are defined as the recurrences within the neighborhood around a given state in phase space. Therefore, persistence ($\theta^{-1}$) estimates the average number of consecutive recurrences, namely the time steps for the system to resemble a given state in its neighboring phase space.
@@
Local dimension $d$ and persistence $\theta^{-1}$ have frequently been associated with the predictability of dynamical systems. More specifically, high local dimensions and low persistence are seen as an indication that the system is less predictable. The rationale for this is as follows. If a system requires more degrees of freedom to be described at a given point in the phase space, then its complexity is greater at that given point, making it possibly less predictable. Similarly, if a system leaves a neighborhood fast, then the dynamics of that neighbour is fast, and non-persistent; aspect that may indicate reduced predictability. While these arguments on predictability are qualitatively valid, and frequently correct, they approach predictability from a complexity ($d$), and mean residence time ($\theta^{-1}$) perspective, without directly addressing predictability.
@@
In this work, we address the predictability of dynamical systems by building upon the framework of local dynamical indices just outlined. In particular, we introduce a new quantity named ‘time-lag recurrence’ and denoted by $\alpha_{\eta}$. This new quantity uses neighboring states to monitor an ensemble of trajectories, and provides an indication in a statistical sense of the local predictability of dynamical systems. This is accomplished in a purely data-driven fashion, i.e. without requiring a model nor further simplifications on the data.
@@
The computation of $\alpha_{\eta}$ is based on the concept of recurrences introduced above, and measured via the logarithmic return quantity defined in equation (1). More specifically, using equation (2), we identify the set of states – i.e., recurrences – $R_{t_{\boldsymbol{\zeta}}}$ in the neighborhood of the reference state $\boldsymbol{\zeta}$ that have positive exceedance, that is:
@@
where $t_{\boldsymbol{\zeta}}$ is the time that identifies the reference state $\boldsymbol{\zeta}$, $w$ is the Theiler window [57] that excludes temporal neighbours of $\boldsymbol{\zeta}$, $N_{n}$ is the number of exceedances (i.e., the number of temporally non-adjacent neighbors) at time $t_{\boldsymbol{\zeta}}$, $t_{u_{i}}$ is the time corresponding to the exceedance $u_{i}$.
After having identified the recurrences $R_{t_{\boldsymbol{\zeta}}}$, to characterise the local predictability of $\boldsymbol{\zeta}$ at time $t_{\boldsymbol{\zeta}}+\eta$ we first need to estimate how neighboring states evolve after time $\eta$ has passed from $t_{\boldsymbol{\zeta}}$. To this end, we first evolve the states within $R_{t_{\boldsymbol{\zeta}}}$ forward by time $\eta$ (where $\eta$ is also referred to as forecasting horizon), which leads to a set of new states, that we name forward recurrences
@@
These new states, that were evolved forward in time by $\eta$ from time $t_{\boldsymbol{\zeta}}$, are represented by the empty blue circles in Fig. 1, and are denoted by using the superscript $(\cdot)^{\eta}$.
Once we have defined the forward recurrences, equation (5), we need to define the forward reference state, that is: the reference state $\boldsymbol{\zeta}={\bf{x}}(t_{\boldsymbol{\zeta}})$ after time $\eta$, and denoted by $\boldsymbol{\zeta}^{\eta}={\bf{x}}(t_{\boldsymbol{\zeta}}+\eta)$. This forward reference state will have its own excedeences ${\bf{u}}({\boldsymbol{\zeta}}^{\eta})$ made of $N_{n}$ elements $\{u_{i}(\boldsymbol{\zeta}^{\eta})\}_{i=1}^{N_{n}}$, and we can compute the forward-reference-state recurrences
@@
in analogy to what was done for the reference state $\boldsymbol{\zeta}$ in equation (4).
With the three sets just introduced in equations (46), namely the recurrences $R_{t_{\boldsymbol{\zeta}}}$ in equation (4), the forward recurrences $R_{t_{\boldsymbol{\zeta}}}^{\eta}$ in equation (5), and the forward-reference-state recurrences $R_{t_{\boldsymbol{\zeta}}+\eta}$ in equation (6), we can define the local predictability index as follows:
@@
where |()||($\cdot$)|| ( ⋅ ) | represents the number of elements within the set ()($\cdot$)( ⋅ ), and $\cap$ is the intersection of two sets. Since both $R_{t_{\boldsymbol{\zeta}}}$ and $R_{{t_{\boldsymbol{\zeta}}+\eta}}$ are defined using the same quantile-based thresholds, the number of elements in one set is equal to the number of elements in the other set, that is: $|R_{{t_{\boldsymbol{\zeta}}}}|=|R_{t_{\boldsymbol{\zeta}}+\eta}|=N_{n}$. In addition, given that $R_{t_{\boldsymbol{\zeta}}}^{\eta}$ is defined from $R_{t_{\boldsymbol{\zeta}}}$, the number of elements it contains is also equal to $N_{n}$. It follows that $0\leq\alpha_{\eta}({\bf{x}}(t))\leq 1$. A value of $\alpha_{\eta}(\boldsymbol{\zeta})$ close to 1 indicates high probability for a neighbor state to remain close to the reference state. We interpret this as a high predictability, in a statistical sense, for state $\boldsymbol{\zeta}$ at time $t_{\boldsymbol{\zeta}}$, and forecasting horizon $\eta$, suggesting that most of its neighboring states stay close to the future trajectory of $\boldsymbol{\zeta}$ as time evolves. Conversely, when $\alpha_{\eta}(\boldsymbol{\zeta})$, is close to zero, this implies that most of the initial neighbors have moved away from the trajectory of $\boldsymbol{\zeta}$ after time $\eta$, indicating low predictability in a statistical sense.
Figure 1: Schematic illustration of the computation of $\alpha_{\eta}(t).$
The derivation of the new quantity $\alpha_{\eta}(\boldsymbol{\zeta})$ relies on the quantile $q$, which is used to define recurrences, and on the distance function $\operatorname{dist}$. The selection of the quantile $q$ can be adapted for different application scenarios, and it reflects the local statistical predictability for different scales. This is not surprising as we are defining $\alpha$ using extreme value theory where quantile choice is always an important step. In Supplementary Information Section B, we show an example of this scale-dependent behaviour using various $q$. This dependence on $q$ can be seen as a favorable properties as it allows for analyzing predictability at different scales. Similarly, the selection of the distance function $\operatorname{dist}$ can be adapted depending on the application, noting that for physical systems, Euclidean distance has been shown to be an appropriate choice [47].
In addition, the new quantity $\alpha_{\eta}(\boldsymbol{\zeta})$ is inherently connected to information theory, since estimating local predictability can be viewed as a process that quantifies the uncertainty or randomness within the system. In fact, if we take a conditional probability perspective, where $P(A|B)$ represents the conditional probability of $A$ given $B$, then we can reformulate the definition of $\alpha_{\eta}$ in equation 7 as follows:
@@
where $A={\bf{x}}(t+\eta)\in R_{{t_{\boldsymbol{\zeta}}+\eta}}$ corresponds to recurrences of the reference state $\boldsymbol{\zeta}$ that, after being propagated forward in time by $\eta$, belong to the forward-state neighborhood $\boldsymbol{\zeta}^{\eta}$ (i.e., they are forward-reference-state recurrences), and $B={\bf{x}}(t)\in R_{t_{\boldsymbol{\zeta}}}$ corresponds to the recurrences of the reference state $\boldsymbol{\zeta}$.
@@
It is helpful here to emphasize some key differences to other data-driven ways of measuring local predictability properties. The classical algorithms proposed in [58] to measure the Lyapunov exponents from available time series allow to estimate the growth rate of the $n-$dimensional volumes (for $n$ smaller than of equal to the dimension of the phase space) defined by $n+1$ initially nearby points. The strategy there strongly relies on using delay coordinates in order to take advantage of Takens’ theorem [59] and reconstruct the key properties of the attractor. By selecting nearby points belonging to a region of interest, local properties can be extracted. Following such geometrical construction, one obtains that
$P[\mathrm{dist}(\bf{x}(t+\eta),\zeta^{\eta})<\epsilon|\mathrm{dist}(\bf{x}(t),\zeta)<\epsilon]\approx e^{-({\sum}_{\lambda_{j}>0}\lambda_{j}^{\zeta})\eta}.$ (9)
where the exponent is proportional to sum of positive local, finite-time Lyapunov exponents. Clearly $\epsilon$ has to be small. After comparing Eqs. 8 and 9 it is tempting to deduce $\alpha_{\eta}\approx e^{-({\sum}_{\lambda_{j}>0}\lambda_{j}^{\zeta})\eta}$ (see later discussion on specific numerical examples). Yet, a key difference between the two definitions exist when one applied them to actual data, because in Eq. 8 one considers neighbourhoods of $\zeta$ and $\zeta^{\eta}$ having the same number of points (thanks to the use of quantiles in defining the threshold of the distance observables), so that such neighbourhoods will in general have a different radius (which is instead identical and equal to $\epsilon$ in Eq. 9).
@@
We have computed the new metric $\alpha_{\eta}(\boldsymbol{\zeta})$ for both idealized and real-world dynamical systems of varying complexity, including the Lorenz-63 attractor and atmospheric circulation patterns. These results are reported in the following subsections. Additionally, in the Supplementary Information, we provide several supplementary results for other systems. Following the discussion of results obtained from various systems, we propose a method for utilizing this index in real-time, addressing the issue introduced at the end of the previous section.
Idealized dynamical systems
[…]
Figure 2: Distribution of $\alpha_{\eta}$ at different prediction horizons for the Lorenz-63 system. (a-e) Lorenz attractor colored by $\alpha_{\eta}$ values at five different forcasting horizons: $\eta=[0.05,0.1,0.2,1,2]\eta_{\ell}$, each corresponding to timesteps: $L=[11,22,44,220,440]$. (f) Probability distribution of $\alpha_{\eta}$ at the same five different forecasting horizons.
Figs. 2(a-e) depict the Lorenz attractor in the phase space colored by the values of $\alpha_{\eta}$, for the 5 different forecasting horizons considered. For each panel in Fig. 2(a-e), the title on the top reports the forecasting horizon, and the average of the local predictability, namely $\bar{\alpha}_{\eta}$. We note how, as we increase the forecasting horizon, $\bar{\alpha}_{\eta}$ tends to decrease, as also depicted by the increasingly darker colors from Fig. 2(a) to Fig. 2(e). This result is in agreement with what we shall expect – the longer the forecasting horizon, the less predictable the system. This conclusion is also summarized in Fig. 2(f), that shows the distribution of $\alpha_{\eta}$ for the five different forecasting horizons: the two longest horizons display a distribution peaked on low values of predictability (<0.2), while shorter horizons tend to display distributions peaked toward higher values of predictability.

### v2

# Time-Lagged Recurrence: a data-driven method to estimate the predictability of dynamical systems

@@ Abstract
Nonlinear dynamical systems are ubiquitous in nature and they are hard to forecast. Not only they may be sensitive to small perturbations in their initial conditions, but they are often composed of processes acting at multiple scales. Classical approaches based on the Lyapunov spectrum rely on the knowledge of the dynamic forward operator, or of a data-derived approximation of it. This operator is typically unknown, or the data are too noisy to derive its faithful representation. Here we propose a new data-driven approach to analyze the local predictability of dynamical systems. This method, based on the concept of recurrence, is closely linked to the well-established framework of local dynamical indices. When applied to both idealized systems and real-world datasets arising from large-scale atmospheric fields, our new approach proves its effectiveness in estimating local predictability. Additionally, we discuss its relationship with other local dynamical indices, and how it reveals the scale-dependent nature of predictability. Furthermore, we explore its link to information theory, its extension that includes a weighting strategy, and its real-time application. We believe these aspects collectively demonstrate its potential as a powerful diagnostic tool for complex systems.
@@ 1 Introduction
The origin of this practical unpredictability may stem either from an intrinsic stochastic nature of the system or from the ignorance of the current state of the system and/or of the dynamic rule controlling its evolution [23], or from a combination of both. Furthermore, it is well-known that even low-dimensional nonlinear systems may be difficult to forecast if they exhibit an exponential divergence of close-by states [19]. In such cases the unpredictability comes from our lack of precise knowledge of the current state of the system: if we knew exactly the initial conditions we would always get the same future evolution. A common quantity used to characterize the predictability of a system is the maximum Lyapunov exponent $\lambda_{\rm{max}}$, which measures the fastest possible rate of exponential divergence between two nearby trajectories in the phase space for an infinitely long time horizon. If we look at the evolution of two trajectories starting from two points close to each other, namely ${\bf{x}}(t_{0})$ and ${\bf{x}}(t_{0})+{\delta}{\bf{x}}$, these will diverge exponentially fast as time progresses, until their distance will plateau around a value bounded by the diameter of the attractor [24]. Even more instructive is to look at the evolution of a ball around the state ${\bf{x}}(t)$. From this, one can define a full Lyapunov spectrum, made of Lyapunov exponents equal in number to the variables used to describe the phase space. The Lyapunov exponents are mean logarithmic growth rates of the lengths of the principal axes of this ball evolving under the action of the dynamic rule [25, 4, 24, 5]. The existence of at least one positive Lyapunov exponent is the hallmark of chaotic dynamical systems. Therefore, measuring predictability in dynamical systems can be re-framed as computing the divergence of a system’s nearby trajectories over time.
@@
Besides Lyapunov exponents and their variants, information theory also offers methods to quantify predictability [39, 23]. Within this framework, various entropy-based indicators are usually employed as predictability measures for model-based prediction, with varying degrees of success [40, 41, 42, 43]. However, because these methods are model-based, they are unsuitable for observational datasets. Other indices from nonlinear time series analysis also use predictability measures such as sample entropy, permutation entropy, and others [44, 45, 46]; nevertheless, they are only effective with time series or systems with relatively low dimensionality. To the best of the authors’ knowledge, no information theory method can infer local predictability information from high-dimensional observational data.
From the brief review of relevant literature just outlined, the above mentioned indices have several drawbacks. More specifically, Lyapunov exponents and their variants may not work well for nonlinear regimes, while information-theory-based methods may not be directly applicable to high-dimensional dynamical systems. In the next section, we leverage the recent framework proposed by Lucarini et al. [47], which uses recurrences to characterize local dynamical properties, to develop a novel local predictability index that differs from existing metrics, such as the Lyapunov exponents and their variants.
@@
Lucarini et al., [47] introduced a novel framework to measure the local properties of dynamical systems. In particular, they combine extreme value theory [48] and the Poincaré recurrence theorem [2] to construct purely data-driven local dynamical indices, that can measure the local dimension of dynamical systems and their persistence at any point in time. Under the ergodic assumption, the Poincaré recurrence theorem guarantees that the system will visit an arbitrarily small neighborhood of any state again and again as time passes. In this framework, for one given state of interest ${\boldsymbol{\zeta}}={\bf{x}}(t={t}_{\boldsymbol{\zeta}})$ in the phase space, its neighboring states, also known as recurrences, are used to characterize its local dynamical properties. This can be formalized by defining the negative logarithmic distance between $\boldsymbol{\zeta}$} and all other states
@@
where the dist function can be any distance metric, and by subsequently defining a quantity called exceedance for the neighboring states (i.e., recurrences):
${\bf{u}}({\boldsymbol{\zeta}})=g({\boldsymbol{\zeta}},{\bf{x}}(t))-s(q,{\boldsymbol{\zeta}}),\;\;\;\;\forall~g({\boldsymbol{\zeta}},{\bf{x}}(t))>s(q,{\boldsymbol{\zeta}}).$ (2)
@@
In Eq. 2, $s(q,{\boldsymbol{\zeta}})$ is a high threshold corresponding to the $q$-quantile of the series $g\left({\boldsymbol{\zeta}},{\bf{x}}(t)\right)$, noting that having $g({\boldsymbol{\zeta}},{\bf{x}}(t))$ above the threshold $s(q,{\boldsymbol{\zeta}})$ is a synonym of requiring a trajectory to fall within a neighborhood of $\boldsymbol{\zeta}$} (intended as a ball).
@@
Building on the above-defined quantities and leveraging extreme value theory, this framework introduces two local dynamical indices: the local dimension $d$ and persistence $\Theta$ (see Materials and Methods for a detailed explanation and derivation), both of which are frequently associated with the predictability of dynamical systems [49, 50, 51]. More specifically, high local dimension and low persistence are seen as an indication that the system is less predictable. The rationale for this is as follows. If a system requires more degrees of freedom (high $d$) to be described at a given point in the phase space, then its complexity is greater at that given point, making it possibly less predictable. Similarly, if a system leaves a neighborhood quickly (low $\Theta$), then the dynamics in that neighborhood are fast and non-persistent, an aspect that may indicate reduced predictability. While these arguments on predictability are qualitatively valid, and frequently correct, they approach predictability from a complexity ($d$), and mean residence time ($\Theta$) perspective, without directly addressing predictability.
Figure 1: Schematic illustration of the computation of $\alpha_{\eta}(t)$, demonstrated in the phase space of the Lorenz-63 system. This figure presents all steps involved in computing the time-lagged recurrence for the reference state $\zeta$ at a forecasting horizon $\eta$, namely $\alpha_{\eta}(\zeta)$. The panels in the second row provide a zoomed-in view of the phase space region where $\alpha_{\eta}(\zeta)$ is defined. Recurrences ($R_{t_{\zeta}}$), forward recurrences ($R_{t_{\zeta}}^{\eta}$), and forward-reference-state recurrences ($R_{t_{\zeta}+{\eta}}$) are represented by solid blue dots, empty blue dots, and orange dots, respectively. The blue circle with radius $e^{-s(q,\zeta)}$ indicates the hypersphere used to define the neighborhood of the reference state, while the orange circle with radius $e^{-s(q,\zeta^{\eta})}$ corresponds to the forward-reference-state.
@@
In this work, we address the predictability of dynamical systems by building upon the framework of local dynamical indices just outlined. In particular, we introduce a new quantity named ‘time-lagged recurrence’ and denoted by $\alpha_{\eta}$. This new quantity uses neighboring states to monitor an ensemble of trajectories, and provides an indication in a statistical sense of the local predictability of dynamical systems. This is accomplished in a purely data-driven fashion, i.e. without requiring a model nor further simplifications on the data.
@@
The computation of $\alpha_{\eta}$ also relies on the concept of recurrences and the quantities defined above in Eqs. 1 and 2, with a schematic illustration presented in Fig. 1. More specifically, we identify the set of states – i.e., recurrences – $R_{t_{\boldsymbol{\zeta}}}$ in the neighborhood of the reference state $\boldsymbol{\zeta}$, that is:
@@
where $u_{i}(\boldsymbol{\zeta})$ represents the exceedances, defined in Eq. 2, and with positive values indicating proximity to the reference state in phase space, $t_{\boldsymbol{\zeta}}$ is the time that identifies the reference state $\boldsymbol{\zeta}$, $N_{n}$ is the number of recurrences (i.e., the number of temporally non-adjacent neighbors) at time $t_{\boldsymbol{\zeta}}$, $t_{u_{i}}$ is the time corresponding to the exceedance $u_{i}$, and $w$ is the Theiler window [52] that excludes temporal neighbors of $\boldsymbol{\zeta}$. We note that if $w$ is chosen too small that cannot exclude temporal neighbors it will lead to spurious high values of $\alpha_{\eta}$.
After having identified the recurrences $R_{t_{\boldsymbol{\zeta}}}$, to characterise the local predictability of $\boldsymbol{\zeta}$, we first need to estimate how neighboring states evolve after time $\eta$ has passed from $t_{\boldsymbol{\zeta}}$. To this end, we first evolve the states within $R_{t_{\boldsymbol{\zeta}}}$ forward by time $\eta$ (where $\eta$ is also referred to as forecasting horizon), which leads to a set of new states, that we name forward recurrences
@@
These new states, that were evolved forward in time by $\eta$ from time $t_{\boldsymbol{\zeta}}$, are represented by the circles with blue contours in Fig. 1, and are denoted by using the superscript $(\cdot)^{\eta}$.
Once the forward recurrences have been defined (Eq. 4), we need to define the forward reference state, that is: the reference state $\boldsymbol{\zeta}={\bf{x}}(t_{\boldsymbol{\zeta}})$ after time $\eta$, and denoted by $\boldsymbol{\zeta}^{\eta}={\bf{x}}(t_{\boldsymbol{\zeta}}+\eta)$. This forward reference state will have its own excedeences ${\bf{u}}({\boldsymbol{\zeta}}^{\eta})$ made of $N_{n}$ elements $\{u_{i}(\boldsymbol{\zeta}^{\eta})\}_{i=1}^{N_{n}}$, and we can compute the forward-reference-state recurrences
@@
in analogy to what was done for the reference state $\boldsymbol{\zeta}$ in Eq. 3.
With the three sets just introduced in Eqs. 35, namely the recurrences $R_{t_{\boldsymbol{\zeta}}}$ in Eq. 3, the forward recurrences $R_{t_{\boldsymbol{\zeta}}}^{\eta}$ in Eq. 4, and the forward-reference-state recurrences $R_{t_{\boldsymbol{\zeta}}+\eta}$ in Eq. 5, we can define the local predictability index as follows:
@@
where $|(\cdot)|$ represents the number of elements within the set $(\cdot)$, and $\cap$ is the intersection of two sets. Since both $R_{t_{\boldsymbol{\zeta}}}$ and $R_{{t_{\boldsymbol{\zeta}}+\eta}}$ are defined using the same quantile-based thresholds, the number of elements in one set is equal to the number of elements in the other set, that is: $|R_{{t_{\boldsymbol{\zeta}}}}|=|R_{t_{\boldsymbol{\zeta}}+\eta}|=N_{n}$. In addition, given that $R_{t_{\boldsymbol{\zeta}}}^{\eta}$ is defined from $R_{t_{\boldsymbol{\zeta}}}$, the number of elements it contains is also equal to $N_{n}$. It follows that $0\leq\alpha_{\eta}({\bf{x}}(t))\leq 1$. A value of $\alpha_{\eta}(\boldsymbol{\zeta})$ close to 1 indicates high probability for a neighbor state to remain close to the reference state. We interpret this as a high predictability, in a statistical sense, for state $\boldsymbol{\zeta}$ at time $t_{\boldsymbol{\zeta}}$, and forecasting horizon $\eta$, suggesting that most of its neighboring states stay close to the future trajectory of $\boldsymbol{\zeta}$ as time evolves. Conversely, when $\alpha_{\eta}(\boldsymbol{\zeta})$, is close to zero, this implies that most of the initial neighbors have moved away from the trajectory of $\boldsymbol{\zeta}$ after time $\eta$, indicating low predictability in a statistical sense.
The derivation of the new quantity $\alpha_{\eta}(\boldsymbol{\zeta})$ relies on the distance function $\operatorname{dist}$ and the quantile $q$, both of which are used to define recurrences. Although the Euclidean distance has been shown to be an appropriate choice for physical systems [47], the distance function can be adjusted depending on the prior knowledge of the system and the application context. The selection of the quantile $q$ can be adapted for different application scenarios, and it reflects the local statistical predictability for different scales. This is not surprising as we are defining $\alpha_{\eta}$ using extreme value theory where quantile choice is always an important step. In SI Appendix section 2, we show an example of this scale-dependent behavior using various $q$. This dependence on $q$ can be seen as a favorable property as it allows for analyzing predictability at different scales.
In addition, the new quantity $\alpha_{\eta}(\boldsymbol{\zeta})$ is inherently connected to information theory, since estimating local predictability can be viewed as a process that quantifies the uncertainty or randomness within the system. In fact, if we take a conditional probability perspective, where $P(A|B)$ represents the conditional probability of $A$ given $B$, then we can reformulate the definition of $\alpha_{\eta}$ in Eq. 6 as follows:
@@
where $A={\bf{x}}(t+\eta)\in R_{{t_{\boldsymbol{\zeta}}+\eta}}$ corresponds to recurrences of the reference state $\boldsymbol{\zeta}$ that, after being propagated forward in time by $\eta$, belong to the forward-state neighborhood $\boldsymbol{\zeta}^{\eta}$ (i.e., they are forward-reference-state recurrences), and $B={\bf{x}}(t)\in R_{t_{\boldsymbol{\zeta}}}$ corresponds to the recurrences of the reference state $\boldsymbol{\zeta}$. We can further link $\alpha_{\eta}$ to information theory using the Shannon entropy, when defined – see SI Appendix section 3.
@@
It is helpful here to emphasize some key differences to other data-driven ways of measuring local predictability properties. The classical algorithms proposed in [53] to measure the Lyapunov exponents from available time series allow to estimate the growth rate of the $n$-dimensional volumes (for $n$ smaller than or equal to the dimension of the phase space) defined by $n+1$ initially nearby points. The strategy there strongly relies on using delay coordinates in order to take advantage of Takens’ theorem [54] and reconstruct the key properties of the attractor. By selecting nearby points belonging to a region of interest, local properties can be extracted. Following such geometrical construction, one obtains that
$P[\mathrm{dist}(\bf{x}(t+\eta),\zeta^{\eta})<\epsilon|\mathrm{dist}(\bf{x}(t),\zeta)<\epsilon]\approx e^{-\left({\sum}_{\lambda_{j}>0}\lambda_{j}^{\zeta}\right)\eta}$ (8)
where the exponent is proportional to the sum of positive local, finite-time, Lyapunov exponents. Clearly $\epsilon$ has to be small. After comparing Eqs. 7 and 8 it is tempting to deduce $\alpha_{\eta}\approx e^{-({\sum}_{\lambda_{j}>0}\lambda_{j}^{\zeta})\eta}$ (see later discussion on specific numerical examples). Yet, a key difference between the two definitions exist when one applies them to actual data, because in Eq. 7 one considers neighborhoods of $\zeta$ and $\zeta^{\eta}$ having the same number of points (thanks to the use of quantiles in defining the threshold of the distance observables), so that such neighborhoods will in general have a different radius (which is instead identical and equal to $\epsilon$ in Eq. 8).
@@
We have computed the new metric $\alpha_{\eta}(\boldsymbol{\zeta})$ for systems of varying complexity to show its effectiveness. The following subsections provide a detailed discussion of the results for the Lorenz-63 attractor and climate data characterizing atmospheric circulation in the Euro-Atlantic sector. Additionally, a brief summary of $\alpha_{\eta}(\boldsymbol{\zeta})$ applied to other systems is presented in a subsequent section to illustrate its broad applicability. Following the discussion of results obtained from various systems, we propose a method for utilizing this index in real-time, addressing the issue introduced at the end of the previous section.
Lorenz-63
We consider the Lorenz-63 attractor as an example of idealized dynamical system. The data for this test case is obtained from the numerical simulation of the Lorenz-63 system, with a time resolution of $5\text{\times}{10}^{-3}$ over $1\text{\times}{10}^{5}$ time steps, using the classical set of parameters that leads the system to be chaotic (see SI Appendix section 1 for details). Figure 2 shows the distribution of $\alpha_{\eta}$ for the Lorenz-63 system, for five different forecasting horizons: $\eta=[0.05,0.1,0.2,1,2]\eta_{\ell}$. The quantity $\eta_{\ell}$, that represents the Lyapunov time, is equal to $1.1$ time units and is computed as the inverse of the maximum Lyapunov exponent of the Lorenz-63 system.
Figure 2: Distribution of $\alpha_{\eta}$ at different prediction horizons for the Lorenz-63 system. (a-e) Lorenz attractor colored by $\alpha_{\eta}$ values at five different forcasting horizons: $\eta=[0.05,0.1,0.2,1,2]\eta_{\ell}$, each corresponding to timesteps: $L=[11,22,44,220,440]$. (f) Probability distribution of $\alpha_{\eta}$ at the same five different forecasting horizons. The quantile $q$ applied in this analysis is 0.99, and the Theiler window size $w$ is set to 50 time steps.
Figs. 2(a-e) depict the Lorenz attractor in the phase space colored by the values of $\alpha_{\eta}$, for the 5 different forecasting horizons considered. For each panel in Fig. 2(a-e), the title on the top reports the forecasting horizon, and the average of the local predictability, namely $\bar{\alpha}_{\eta}$. We note how, as we increase the forecasting horizon, $\bar{\alpha}_{\eta}$ tends to decrease, as also depicted by the increasingly darker colors from Fig. 2(a) to Fig. 2(e). This result is in agreement with what we shall expect – the longer the forecasting horizon, the less predictable the system. This conclusion is also summarized in Fig. 2(f), that shows the distribution of $\alpha_{\eta}$ for the five different forecasting horizons: the two longest horizons display a distribution peaked on low values of predictability ($\sim 0.2$), while shorter horizons tend to display distributions peaked toward higher values of predictability. To further test the resilience of $\alpha_{\eta}$ to noise, we applied $\alpha_{\eta}$ to the Lorenz-63 system with stochastic diffusion terms for comparison (see SI Appendix Fig. S1). The results are similar to Fig. 2, while we observe small-scale fluctuations of $\alpha_{\eta}$ over the attractor and a more rapid drop in average predictability, as one shall expect.